What is the probability that a hired bike will be deemed beyond repair provided that the wheels get damaged independently of each other



1
Expert's answer
2022-05-10T06:59:22-0400

Let A denote front wheel damaged. So "P(A)=0.08"

Let B denote that back wheels get damaged. So "P(B)=0.12."

As the two events are independent


"P(A\cap B)=P(A)P(B)"

The probability that a hired bike will be deemed beyond repair  is


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=P(A)+P(B)-P(A)P(B)"

"=0.08+0.12-0.08(0.12)=0.1904"